the ratio of the length of a rectangle to its breadth is 9:4. if the perimeterof rectangle is 104cm, find the area of the rectangle.
Answer by oberobic(2304)

The ratio of L:W = 9:4.
L = 9x
W = 4x
.
Perimeter = 104 cm
P = 9x + 4x + 9x + 4x
26x = 104
x = 4
.
L = 9*4 = 36 cm
W = 4*4 = 16 cm
.
Check the perimeter.
.
2(36) + 2(16) = 72 + 32 = 104
.
Re-read the question to determine if you've answered it.
Nope. It asks for the area of the rectangle.
.
Answer: Area = L*W = 36*16 = 576 cm^2
.
Done.